# Large-scale Pre-training for Grounded Video Caption Generation

### v2

@@ Appendix D Details of the automatic annotation method
[…]
@@
Appendix E Protocol for human annotations
In Figure 9, we describe the annotation guidelines for annotating the training/validation/test sets of the GROVE dataset.
@@ Appendix E Protocol for human annotations
Figure 10: The full prompt for Stage 2 (Video-level caption aggregation) of our automatic annotation approach (Section 3 in the main paper).
@@
Figure 11: The full prompt for Stage 3 (Temporally consistent bounding box annotation) of our automatic annotation approach (Section 3 in the main paper).
Appendix F Prompts for automatic curation of spatio-temporally grounded captions
The full prompt for the Stage 2 (Video-level caption aggregation) of our automatic annotation approach (Section 3 in the main paper) is shown in Figure 10 and the full prompt for Stage 3 (Temporally consistent bounding box annotation) in Figure 11.

### v3

@@ Appendix F Details of the automatic annotation method
Additional details of Stage 3. We provide additional details of the procedure of Stage 3 using the example from Figure 2, right. The object in the woman’s hands is described as ‘a green beverage’ and ‘a glass of green liquid’ across different frames. Stage 2 has provided the video-level noun phrases ‘a woman’ and ‘a beverage’. Stage 3 is formulated as a classification problem where each one of ‘a green beverage’ and ‘a glass of green liquid’ are the inputs to be classified in one of the classes {‘a woman’, ‘a beverage’, $\emptyset$} and thus associated with the right bounding box. The class $\emptyset$ represents the “None” class, i.e. when an input does not belong to any of the known classes and it is useful for noisy inputs.
@@
Appendix G Protocol for human annotations
In Figure 10, we describe the annotation guidelines for annotating the training/validation/test sets of the iGround dataset.
@@ Appendix G Protocol for human annotations
Figure 11: The full prompt for Stage 2 (Video-level caption aggregation) of our automatic annotation approach (Section 3).
@@
Figure 12: The full prompt for Stage 3 (Temporally consistent bounding box annotation) of our automatic annotation approach (Section 3).
Appendix H Prompts for automatic curation of spatio-temporally grounded captions
The full prompt for the Stage 2 (Video-level caption aggregation) of our automatic annotation approach (Section 3) is shown in Figure 11 and the full prompt for Stage 3 (Temporally consistent bounding box annotation) in Figure 12.